OpenStudy (anonymous):

Solve for x: 5 over x squared minus 4 plus 2 over x equals 2 over x minus 2

OpenStudy (anonymous):

Do you mean:\[5/x^2-4+2/x=2/x-2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not\[5/(x^2-4)+2/x=2/(x-2)\]

OpenStudy (anonymous):

isnt that technically the same thing ?

OpenStudy (anonymous):